# SuperBPE: Space Travel for Language Models

### v2

@@ 4.2 Fine-grained analysis of bits-per-byte
The reduction in low-loss tokens also makes sense from a qualitative linguistic analysis of SuperBPE tokens. In Table 3, we show the most common POS tags among superword tokens in SuperBPE, along with random examples for each tag. The tokens often capture common multi-word expressions (by accident, of course, for a living) that function as a single semantic unit (Schneider et al., 2014). As an example, prepositions (IN) figure prominently in superword tokens (e.g., depend on, distinction between) and require lexeme-specific memorization. The individual words in these fixed expressions are often semantically vacuous and have little possible variation in context, so they are easy to predict once memorized.
@@
On the other hand, the much thinner tail of very high-loss tokens shows that, in the worst case, the SuperBPE model consistently puts more probability mass on the correct token. On average, we expect models to suffer high loss on tokens that are difficult to predict. This may explain why SuperBPE can outperform BPE on downstream tasks while having higher average BPB: the tokens that are scored in task evaluations tend to be among the hardest to predict. This is consistent with prior findings that models generally continue to improve in downstream tasks even as their overall loss plateaus, due to improving on a narrow and difficult slice of the distribution (Liu et al., 2023).
@@ 4.3 Scaling
To characterize the scaling behavior of SuperBPE, we also perform experiments at smaller scales.999For scaling, we focus on BPB, since our downstream evals are too noisy for our small models to make meaningful comparisons. We train baseline models at 680M and 1.9B, and scale the base number of training tokens proportionately to the number of parameters. We also perform runs at , , and the base number of tokens to observe the trend with respect to training duration. Then, we train two SuperBPE models that match the training budget of each baseline BPE model: one that matches the baseline in parameter count (analogous to SuperBPE 8B), and a larger model that matches in both train and inference compute (analogous to SuperBPE 11B). We focus on the 180k tokenizer to reduce complexity.
We plot BPB at the end of training for each run in Figure 5. In the undertrained regime, both SuperBPE models achieve lower BPB than the baseline. In the overtrained regime, the ranking from worst to best is SuperBPE (matching parameter count), BPE, and SuperBPE (matching inference compute). Additionally, the separation between the models increases with further overtraining. We provide more results and comments on scaling in §B.3.
@@ 5 Related Work
Prior work has explored processing text at multiple levels of granularity (Lai et al., 2021; Zhang et al., 2021) or creating multi-word tokens through frequency-based identification of -grams (Gee et al., 2023; Kumar & Thawani, 2022). However, these were explored in limited experimental contexts (mainly for machine translation) and had mixed effectiveness. Naively disabling pretokenization in BPE has been found to severely hurt model performance (Dagan et al., 2024; Schmidt et al., 2024; Kudo, 2018), though it may be more promising for unigram tokenization (Kudo & Richardson, 2018), as adopted by Jurassic (Lieber et al., 2021) and BloombergGpt (Wu et al., 2023). In concurrent work, Huang et al. (2025) disentangle the input and output vocabulary, expanding only the former to include -gram tokens. Their method requires significant modifications of the LM input component and considers fixed length of -gram tokens.
@@
Multi-token prediction (MTP) equips LMs with some extra parameters to predict multiple tokens in a single time step (Qi et al., 2020; Gloeckle et al., 2024), and was recently adopted by Deepseek-V3, though the MTP module is discarded at inference-time. The effectiveness corroborates that LMs are capable of predicting more than one subword in a forward pass. However, these approaches fix the number of tokens predicted in each time step and require modifications to the architecture and training objective. Nonetheless, the benefits of MTP and superword tokens may be orthogonal.
@@
Some works have explored the possibility of removing tokenization entirely from LMs and directly modeling text as a sequence of bytes (Clark et al., 2022; Xue et al., 2022; Wang et al., 2024). To overcome the increased compute requirement due to expanded sequence lengths, alternative architectures have been proposed that either segment bytes into fixed-length patches (Tay et al., 2022; Yu et al., 2023) or dynamically predict patch boundaries with sub-modules (Nawrot et al., 2023; Pagnoni et al., 2024; Ahia et al., 2024), increasing the complexity of the model.
@@ 6 Conclusion
Although tokenization lies at the foundation of language modeling—acting as the lens through which models view text—the algorithms in use have remained largely unchanged over the past decade. SuperBPE builds on the natural observation that tokens need not be limited to subwords, extending the BPE algorithm to superword tokens. When replacing subword BPE tokenizers with SuperBPE tokenizers in pretraining, we find that language models achieve better performance over a large suite of downstream tasks, while also being substantially more efficient at inference time. These benefits are achieved without modifying the underlying model architecture, making SuperBPE a compelling alternative to BPE that seamlessly integrates with modern language model ecosystems.
@@ Acknowledgments
We would like to thank Alex Fang for pretraining advice, Vivek Ramanujan for helping debug our distributed training setup, Ian Magnusson for helpful comments on LM evaluation, and Zhaofeng Wu, Alexander Fang, and Xiaochuang Han for feedback on drafts. We would also like to thank Luca Soldaini, Gonçalo Faria, Shrimai Prabhumoye, Matt Jordan, Artidoro Pagnoni, Mike Lewis, Doug Downey, Shannon Shen, and the UW NLP community for valuable conversations about this work. Both co-first authors, AL and JH, are supported by the NSF Graduate Research Fellowship Program. JH and SO are supported in part by the Microsoft Grant for Customer Experience Innovation. This work was further supported in part by NSF DMS-2134012, NSF CCF-2019844, ONR N00014-24-1-2207, NSF 2113530, and NVIDIA resources provided through the National AI Research Resource Pilot (NAIRR).
@@ A.1.2 Limit on the size of superword tokens
Even after truncating the longest 1% of documents, we found that SuperBPE tokenizers can still have extremely long tokens consisting of highly duplicated boilerplate text such as the Project Gutenberg license or common internet phrases such as You_are_commenting_using_your. This issue is already present in BPE tokenizers trained on Chinese, which contain sentence-long tokens clearly taken from pornographic content. For instance, tokens in Gpt-4o’s tokenizer include {CJK*}UTF8gbsn最新高清无码 = latest HD uncensored and {CJK*}UTF8gbsn娱乐网址 = entertainment website. To prevent concerns about the tokenizer directly revealing parts of the training data (Hayase et al., 2024), we enforce an upper bound of 4 words in our tokens. Empirically, we found that this had no measurable impact on the encoding efficiency of the tokenizers or the resulting trained LMs.
@@ A.2 Scaling model configurations
A.3 Evaluation Suite
@@
contains simple arithmetic problems (Brown et al., 2020).101010https://huggingface.co/datasets/EleutherAI/arithmetic We use the 2da, 2dm, and 2ds splits for addition, multiplication, and division of (up to) 2-digit numbers.
@@
contains open-ended questions from the “Jeopardy!” quiz show. 111111https://www.kaggle.com/datasets/tunguz/200000-jeopardy-questions
@@ Appendix B Additional Results
B.1 Task evaluation
@@ B.2 BPB evaluation
B.3 Additional scaling experiments
@@
B.3.1 Matching model parameter count
@@
B.3.2 Matching inference compute
@@ C.1 Superword token analysis
How many superword tokens are in SuperBPE tokenizers? While the second stage of the pretokenization curriculum allows learning of superword tokens, subword tokens can still be learned. Shown in 8(a), for transition points 80k, the number of superword tokens is relatively steady around 120k. Past 100k, almost all tokens learned in Stage 2 are superword tokens. 8(b) shows the number of whitespace-delimited words in the superword tokens of SuperBPE with 180k.
@@ C.2 Analysis of token frequencies in encoding
We also analyze token frequency statistics under BPE versus SuperBPE tokenizers. 9(a) shows the relation between token rank (in frequency) and frequency. While tokens in BPE demonstrate a standard Zipfian relation, the slope of SuperBPE curves have a more shallow slope, meaning that the rate of decay in token frequency is smaller. The smaller proportion of tokens with very low counts may reduce prevalence and severity of glitch tokens (Rumbelow & Watkins, 2023; Land & Bartolo, 2024).
9(b) shows the minimum number of tokens from the vocabulary needed to cover any given proportion of data. For BPE, the relation is striking—only 57% of tokens are needed to encode 99% of the data! The remaining tokens make up a long tail of infrequent tokens. In contrast, SuperBPE tokenizers make better use of the vocabulary. For 80k and 180k, this statistic is 90% and 70% of tokens, respectively.
@@ C.3 Distributional Distortion at the Prompt Boundary
Several fixes for this issue have been proposed: at training time, token merges can be randomly dropped (Provilkov et al., 2020; Sims et al., 2025; DeepSeek-AI et al., 2025) to expose LMs to the internal makeup of tokens; at inference time, options include token healing (Lundberg, 2023), algorithmic correction (Phan et al., 2024), and enumeration of all relevant segmentations of the prompt (Vieira et al., 2024). We leave a detailed comparison of these techniques to future work.

### v3

@@ 4.2 Loss distribution analysis
The reduction in low-loss tokens also makes sense from a qualitative linguistic analysis of SuperBPE tokens. In Table 3, we show the most common POS tags among superword tokens in SuperBPE along with random examples for each tag. The tokens often capture common multi-word expressions (by accident, of course, for a living) that function as a single semantic unit (Schneider et al., 2014). As an example, prepositions (IN) figure prominently in superword tokens (e.g., depend on, distinction between) and require lexeme-specific memorization. The individual words in these fixed expressions are often semantically vacuous and have little possible variation in context, so they are easy to predict once memorized.
@@
On the other hand, the much thinner tail of very high-loss tokens shows that, in the worst case, the SuperBPE model consistently puts more probability mass on the correct token. On average, we expect models to suffer high loss on tokens that are difficult to predict. This may explain why SuperBPE can outperform BPE on downstream tasks but have higher average BPB: the tokens scored in task evaluations tend to be among the hardest to predict. This is consistent with prior findings that models generally continue to improve in downstream tasks even as their overall loss plateaus due to improving on a narrow and difficult slice of the distribution (Liu et al., 2023).
@@ 4.3 Scaling
To characterize the scaling behavior of SuperBPE, we also perform experiments at smaller scales.101010For scaling, we focus on BPB since our downstream evaluations are too noisy for our small models to make meaningful comparisons. We train baseline models at 680M and 1.9B and scale the base number of training tokens proportionately to the number of parameters. We also perform runs at , , and the base number of tokens to observe the trend with respect to training duration. Then, we train two SuperBPE models that match the training budget of each baseline BPE model, one that matches the baseline in parameter count (analogous to SuperBPE 8B) and a larger model that matches in both train and inference compute (analogous to SuperBPE 11B). We focus on the 180k tokenizer to reduce complexity.
We plot BPB at the end of training for each run in Figure 5. In the under-trained regime, both SuperBPE models achieve lower BPB than the baseline. In the over-trained regime, the ranking from worst to best is SuperBPE (matching parameter count), BPE, and SuperBPE (matching inference compute). Additionally, the separation between the models increases with further over-training. We provide additional results and comments on scaling in §B.4.
@@ 5 Related Work
Prior work has explored processing text at multiple levels of granularity (Lai et al., 2021; Zhang et al., 2021) or creating multi-word tokens through frequency-based identification of -grams (Gee et al., 2023; Kumar & Thawani, 2022). However, these were explored in limited experimental contexts (mainly for machine translation) and had mixed effectiveness. Naively disabling pretokenization in BPE has been found to severely degrade model performance (Dagan et al., 2024; Schmidt et al., 2024; Kudo, 2018), although this approach may be more promising for unigram tokenization (Kudo & Richardson, 2018), as adopted by Jurassic (Lieber et al., 2021) and BloombergGpt (Wu et al., 2023). In concurrent work, Huang et al. (2025) disentangle input and output vocabularies, expanding only the former to include -gram tokens. Their method requires significant modifications of the LM input component and considers fixed length of -grams.
@@
Multi-token prediction (MTP) equips LMs with some extra parameters to predict multiple tokens in a single time step (Qi et al., 2020; Gloeckle et al., 2024) and was recently adopted by Deepseek-V3, though the MTP module is discarded at inference-time. MTP’s effectiveness corroborates that LMs are capable of predicting more than one subword in a forward pass. However, these approaches fix the number of tokens predicted in each time step and require modifications to the architecture and training objective. We note that the benefits of MTP and superword tokens may be orthogonal.
@@
Some works have explored the possibility of completely removing tokenization from LMs and directly modeling text as a sequence of bytes (Clark et al., 2022; Xue et al., 2022; Wang et al., 2024). To overcome the increased compute requirement due to expanded sequence lengths, alternative architectures have been proposed that either segment bytes into fixed-length patches (Tay et al., 2022; Yu et al., 2023) or dynamically predict patch boundaries with sub-modules (Nawrot et al., 2023; Pagnoni et al., 2024; Ahia et al., 2024; Hwang et al., 2025); these dynamic patches have been qualitatively observed to span multiple words.
Tokenizer transfer
Many methods have been proposed to adapt models after training to use a different tokenizer. These may rely on intervention during pretraining (Chen et al., 2023), continued training for a subset of layers (Marchisio et al., 2023), or leveraging self-distillation (Minixhofer et al., 2025), heuristic, (Minixhofer et al., 2022; Gee et al., 2022; Tran, 2020; Liu et al., 2024b; Dobler & De Melo, 2023), or hypernetwork-based (Minixhofer et al., 2024) initialization of a fresh embedding matrix, optionally followed by fine-tuning. These methods may be used to upgrade existing models to use SuperBPE tokenizers, with the goal of reducing inference cost while maintaining performance. We leave this direction to future work.
@@ 6 Conclusion
Although tokenization lies at the foundation of language modeling, acting as the lens through which models view text, the algorithms in use have remained largely unchanged over the past decade. SuperBPE builds on the observation that tokens need not be limited to subwords, extending the BPE algorithm to superword tokens. When replacing subword BPE tokenizers with SuperBPE tokenizers in pretraining, we find that language models perform better over a large suite of downstream tasks, while also being substantially more efficient at inference time. These benefits are achieved without modifying the underlying model architecture, making SuperBPE a compelling alternative to BPE that seamlessly integrates with modern language model ecosystems.
@@ Acknowledgments
We would like to thank Alex Fang for pretraining advice, Vivek Ramanujan for helping debug our distributed training setup, Ian Magnusson for helpful comments on LM evaluation, and Zhaofeng Wu, Alexander Fang, and Xiaochuang Han for feedback on drafts. We are also grateful to Luca Soldaini, Gonçalo Faria, Shrimai Prabhumoye, Matt Jordan, Artidoro Pagnoni, Mike Lewis, Doug Downey, Shannon Shen, and the UW NLP community for valuable conversations about this work. Both co-first authors, AL and JH, are supported by the NSF Graduate Research Fellowship Program. JH and SO are supported in part by the Microsoft Grant for Customer Experience Innovation. This work was partially funded by NSF DMS-2134012, NSF CCF-2019844, ONR N00014-24-1-2207, and NSF 2113530 as well as with NVIDIA resources provided through the National AI Research Resource Pilot (NAIRR).
@@ A.1.2 Limit on the size of superword tokens
Even after truncating the longest 1% of documents, we found that SuperBPE tokenizers can still have extremely long tokens consisting of highly duplicated boilerplate text such as the Project Gutenberg license or common internet phrases such as You_are_commenting_using_your. This issue is already present in BPE tokenizers trained on Chinese, which contain sentence-long tokens clearly taken from pornographic content. For instance, tokens in Gpt-4o’s tokenizer include 最新高清无码 = latest HD uncensored and 娱乐网址 = entertainment website. To prevent concerns about the tokenizer directly revealing parts of the training data (Hayase et al., 2024), we enforce an upper bound of 4 words in our tokens. Empirically, we found that this had no measurable impact on the encoding efficiency of the tokenizers or the resulting trained LMs.
@@ A.2 Scaling model configurations
A.3 Compute used for model training
All models were pretrained on 32 8H100 nodes.
A.4 Evaluation Suite
@@
contains simple arithmetic problems (Brown et al., 2020).111111https://huggingface.co/datasets/EleutherAI/arithmetic We use the 2da, 2dm, and 2ds splits for addition, multiplication, and division of (up to) 2-digit numbers.
@@
contains open-ended questions from the “Jeopardy!” quiz show.121212https://www.kaggle.com/datasets/tunguz/200000-jeopardy-questions
@@ Appendix B Additional Results
B.1 How BPB varies with context length
In our main experiments (§3), we adjust the context size of SuperBPE models to match the effective context size of the BPE model in raw text. To justify this design choice, we show that the next token becomes easier to predict as a function of the preceding context in bytes (not tokens). Figure 6 shows the average BPB at every token index (left) vs byte index (right) — when measured at fixed token indices, SuperBPE has an advantage from seeing more context (achieving lower loss on average at the same token index), whereas at fixed byte indices, this advantage goes away.
B.2 Task evaluation
@@ B.3 BPB evaluation
B.4 Additional scaling experiments
@@
B.4.1 Matching model parameter count
@@
B.4.2 Matching inference compute
@@ C.1 Superword token analysis
How many superword tokens are in SuperBPE tokenizers? While the second stage of the pretokenization curriculum allows learning of superword tokens, subword tokens can still be learned. Shown in 9(a), for transition points 80k, the number of superword tokens is relatively steady around 120k. Past 100k, almost all tokens learned in Stage 2 are superword tokens. 9(b) shows the number of whitespace-delimited words in the superword tokens of SuperBPE with 180k.
@@ C.2 Analysis of token frequencies in encoding
We also analyze token frequency statistics under BPE versus SuperBPE tokenizers. 10(a) shows the relation between token rank (in frequency) and frequency. While tokens in BPE demonstrate a standard Zipfian relation, the slope of SuperBPE curves have a more shallow slope, meaning that the rate of decay in token frequency is smaller. The smaller proportion of tokens with very low counts may reduce prevalence and severity of glitch tokens (Rumbelow & Watkins, 2023; Land & Bartolo, 2024).
10(b) shows the minimum number of tokens from the vocabulary needed to cover any given proportion of data. For BPE, the relation is striking—only 57% of tokens are needed to encode 99% of the data! The remaining tokens make up a long tail of infrequent tokens. In contrast, SuperBPE tokenizers make better use of the vocabulary. For 80k and 180k, this statistic is 90% and 70% of tokens, respectively.
@@ C.3 Distributional Distortion at the Prompt Boundary
Several fixes for this issue have been proposed: at training time, token merges can be randomly dropped (Provilkov et al., 2020; Sims et al., 2025; DeepSeek-AI, 2025) to expose LMs to the internal makeup of tokens; at inference time, options include token healing (Lundberg, 2023), algorithmic correction (Phan et al., 2024), and enumeration of all relevant segmentations of the prompt (Vieira et al., 2024). We leave a detailed comparison of these techniques to future work.